The city of Calgary is prone to severe hailstorms. Insurance agents claim that a home owner in Calgary can expect to replace their rook, due to hail damage, once every 10 years. What is the probability that in 12 years a home owner in Calgary will need to replace their roof twice because of hail?
Answer #1

This is Poisson's distribution, lamda = 1 per 10 years
Here, λ = 12/10 = 1.2 and x = 2
As per Poisson's distribution formula P(X = x) = λ^x * e^(-λ)/x!

We need to calculate P(X = 2)
P(X = 2) = 1.2^2 * e^-1.2/2!
P(X = 2) = 0.2169
Ans: 0.2169